### v2

# Distributed Koopman Learning using Partial Trajectories for Control

Abstract
This paper proposes a distributed data-driven framework for dynamics learning, termed distributed deep Koopman learning using partial trajectories (DDKL-PT). In this framework, each agent in a multi-agent system is assigned a partial trajectory offline and locally approximates the unknown dynamics using a deep neural network within the Koopman operator framework. By exchanging local estimated dynamics rather than training data, agents achieve consensus on a global dynamics model without sharing their private training trajectories. Simulation studies on a surface vehicle demonstrate that DDKL-PT attains consensus with respect to the learned dynamics, with each agent achieving reasonably small approximation errors over the testing data. Furthermore, a model predictive control scheme is developed by integrating the learned Koopman dynamics with known kinematic relations. Results on goal-tracking and station-keeping tasks support that the distributedly learned dynamics are sufficiently accurate for model-based optimal control.
I Introduction
Learning system dynamics from system states-inputs data pairs has attracted considerable research attention due to the increasing complexity of the autonomous systems [1, 2, 3, 4]. In the field of dynamics learning, recent methods include using deep neural networks (DNNs) [5, 6], Physics informed neural networks [7], and lifting linearization methods such as Koopman operator methods [2, 3, 4] have been proven to be effective in approximating unknown dynamical systems.
Among the various approaches, the Koopman operator has gained popularity due to its capacity to represent unknown dynamics in a linear form. Two prominent Koopman-based methods, dynamic mode decomposition [8] and extending dynamic mode decomposition have been proposed to lift the state space to a higher-dimensional space, for which the evolution is approximately linear [9]. However, the problem of selecting basis functions to ensure that the lifted system is linear and accurate remains a significant challenge. To tackle this problem, various eigen decomposition-based truncation techniques have been proposed. For instance, [10] proposed to use deep learning methods to discover the eigenfunctions of the approximated Koopman operator, and [11, 12, 13, 14] employed DNNs as observable functions of the Koopman operator, which are tuned based on collected state-control pairs by minimizing a suitably defined loss function which is also referred as the deep Koopman operator method (DKO). Recent work, such as [15], has extended the DKO method to approximate nonlinear time-varying systems.
Although Koopman-based methods have been successfully applied to dynamics learning, existing approaches do not directly address scenarios involving large-scale datasets of state–input pairs. To overcome this limitation, we propose a consensus-based framework, termed distributed deep Koopman learning using partial trajectories (DDKL-PT), for approximating nonlinear time-invariant systems (NTIS) with linear models to enable applications such as optimal control design. Unlike centralized block-wise Koopman operator methods [16, 17], which identify a single global operator for all agents in a multi-agent system (MAS), DDKL-PT assigns each agent an offline partial trajectory and enables local dynamics estimation via a deep neural network embedded in the Koopman operator framework. Through local communication of the estimated dynamics within neighbors, agents reach consensus on a global Koopman representation without sharing their private trajectory data. This distributed approach (i) reduces the computational burden of large-scale learning by distributing the workload across agents, thereby improving scalability, and (ii) preserves data locality and privacy when trajectories are inherently distributed. The main contributions are summarized as follows:
  • We develop a distributed deep Koopman learning algorithm for identifying the dynamics of unknown NTIS when each agent in a MAS has access only to partial state–input trajectories. The algorithm ensures consensus across agents on the approximated dynamics without sharing private training data during the learning process.
  • We present a model predictive control (MPC) scheme for a surface vehicle performing goal-tracking and station-keeping tasks. The controller integrates the learned Koopman dynamics with known kinematic relations, and simulation results demonstrate that the distributedly learned dynamics are sufficiently accurate for model-based optimal control.
This paper is structured as follows: In Section II, we introduce the problem. Section III details the proposed distributed algorithm. Section IV presents numerical simulations. Finally, Section V concludes with a summary of the results.
Notations. For a matrix $\in\mathbb{R}^{n\times m}, A\parallel A\parallel$ denotes its Euclidean norm, $\parallel A\parallel_{F}$ denotes its Frobenius norm, $A^{\prime}$ denotes its transpose, and $A^{\dagger}$ denotes its Moore-Penrose pseudoinverse. ${\mathbf{I}}_{n}$ denotes a $n\times n$} identity matrix.
II The Problem
Consider the following discrete-time nonlinear time-invariant systems (NTIS):
$\boldsymbol{x}(t+1)=\boldsymbol{f}(\boldsymbol{x}(t),\boldsymbol{u}(t)),$ (1)
where $t=0,1,2,\cdots$ denotes the time index, $\boldsymbol{x}(t)\in\mathbb{R}^{n}$ and $\boldsymbol{u}(t)\in\mathbb{R}^{m}$ denote the system state and control input at time $t$, respectively, and $\boldsymbol{f}:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is the time-invariant mapping function which is assumed to be unknown. A trajectory of system states-inputs from time 0 to time $T$ is represented by:
$\boldsymbol{\xi}=\{(\boldsymbol{x}_{t},\boldsymbol{u}_{t}):t=0,1,2,\cdots,T\},$ (2)
where $(\boldsymbol{x}_{t},\boldsymbol{u}_{t})$ denotes the observed fixed state–input pair, which is distinguished from the state-input variables $(\boldsymbol{x}(t),\boldsymbol{u}(t))$ used elsewhere in this paper.
The goal of the centralized deep Koopman operator (DKO) is to, given the entire trajectory $\boldsymbol{\xi}$, identify constant matrices $A^{*}\in\mathbb{R}^{r\times r}$, $B^{*}\in\mathbb{R}^{r\times m}$, $C^{*}\in\mathbb{R}^{n\times r}$, and a parameter vector $\boldsymbol{\theta}^{*}\in\mathbb{R}^{p}$ such that for any $0\leq t\leq T-1$, the following holds:
$\boldsymbol{x}_{t+1}=C^{*}\Big(A^{*}\boldsymbol{g}(\boldsymbol{x}_{t},\boldsymbol{\theta}^{*})+B\boldsymbol{u}_{t}\Big),$ (3)
where $\boldsymbol{g}(\cdot,\boldsymbol{\theta}^{*}):\mathbb{R}^{n}\rightarrow\mathbb{R}^{r}$ is a function with known structure, and the lifting dimension satisfies $r\geq n$.
Consider a group of $N\geq 1$ agents with its node set denoted as $\mathcal{V}=\{1,2,\dots,N\}$. Each agent $i\in\mathcal{V}$ can receive/send information from/to its neighbor set $\mathcal{N}_{i}\subset\mathcal{V}$, $i\in\mathcal{N}_{i}$. In other words, the communication between agent $i$ and agent $i$’s neighbors in $\mathcal{N}_{i}$ is bidirectional. Neighbor relations between distinct pairs of agents are characterized by a self-arced undirected graph $\mathbb{G}=\{\mathcal{V},\mathcal{E}\}$ such that $(i,j)\in\mathcal{E}$ if and only if agents $i$ and $j$ are neighbors. We assume $\mathbb{G}$ is connected. As shown in Fig. 1, we assume a MAS of which any agent $i$ observes partial trajectory of (2) denoted by
$\boldsymbol{\xi}_{i}=\{(x_{t_{i,k}},u_{t_{i,k}}):k=0,1,2,\cdots,T_{i}\},$ (4)
where $t_{i,0}$ and $t_{i,T_{i}}$ denotes the starting time and end time of trajectory $\boldsymbol{\xi}_{i}$ respectively with $0\leq t_{i,0}<t_{i,T_{i}}\leq T$. Note that no restrictions are imposed on $\boldsymbol{\xi}_{i}$; it may contain as few as two data pairs, and trajectory segments across different agents may overlap.
Figure 1: Illustration of the distributed Koopman learning, where each agent is only available for a partial trajectory.
When the dataset $\boldsymbol{\xi}$ contains a large number of state–input pairs (i.e., $T$ is large), a centralized DKO learner may face significant difficulties in efficiently processing the data. Moreover, the partially observed trajectory $\boldsymbol{\xi}_{i}$ available to each agent is generally insufficient for independently identifying the dynamics in (3). Motivated by this limitation, the problem of interest is to obtain (3) by developing a distributed DKO framework, wherein each agent $i$ maintains its own estimated dynamics set
$\mathcal{K}_{i}=\{A_{i},B_{i},C_{i},\boldsymbol{\theta}_{i}\},$ (5)
which serves as a local approximation of $\{A^{*},B^{*},C^{*},\boldsymbol{\theta}^{*}\}$ in (3). Agents iteratively update their estimates $\mathcal{K}_{i}$ via both local partial trajectory and collaboratively exchanging $\mathcal{K}_{i}$ with their neighbors.
III The Algorithm
In this section, we outline the main challenges and key ideas underlying the proposed problem, followed by a proposed distributed algorithm for achieving dynamics in (3).
III-A Challenges and Key ideas
The main challenge in achieving (3) lies in the fact that each agent $i$ in the MAS only observes partial trajectories of (2), which are insufficient to recover (3) for all state-input pairs in (2). To overcome this limitation, we propose a distributed algorithm in which each agent learns its local dynamics from available training data and subsequently exchanges the learned dynamics with its neighbors to estimate (3). This procedure can be formulated as a multi-agent optimization problem.
Local dynamics estimation. Suppose a MAS of $N$ agents, where each agent $i$ can only access a set of trajectory segments $\boldsymbol{\xi}_{i}$ from (4). Following Koopman operator theory [18], to approximate the unknown dynamics in (1), each $i$ is to identify constant matrices $A_{i}^{*}\in\mathbb{R}^{r\times r}$, $B_{i}^{*}\in\mathbb{R}^{r\times m}$, $C_{i}^{*}\in\mathbb{R}^{n\times r}$, and a parameter vector $\boldsymbol{\theta}_{i}^{*}\in\mathbb{R}^{p}$ such that the following holds:
$\boldsymbol{x}_{t_{i,k+1}}=C_{i}^{*}\Big(A_{i}^{*}\boldsymbol{g}(\boldsymbol{x}_{t_{i,k}},\boldsymbol{\theta}_{i}^{*})+B_{i}^{*}\boldsymbol{u}_{t_{i,k}}\Big).$ (6)
To this end, we define for each agent $i$ the local dynamics learning error function over its partially observed trajectory as
[…]
where the first term is designed to approximate a linear mapping between $\boldsymbol{x}(t)$ and $\boldsymbol{g}(\boldsymbol{x}(t),\boldsymbol{\theta}_{i})$, while the second term enforces the approximate evolution from the lifted state $\boldsymbol{g}(\boldsymbol{x}(t),\boldsymbol{\theta}_{i})$ and input $\boldsymbol{u}(t)$ to $\boldsymbol{g}(\boldsymbol{x}(t+1),\boldsymbol{\theta}_{i})$. Furthermore, we introduce the following data matrices built from $\boldsymbol{\xi}_{i}$:
${\mathbf{X}}_{i}$ $=\begin{bmatrix}\boldsymbol{x}_{t_{i,0}},\boldsymbol{x}_{t_{i,1}},\cdots,\boldsymbol{x}_{t_{i,T_{i}-1}}\end{bmatrix}\in\mathbb{R}^{n\times T_{i}},$ (8)
$\bar{{\mathbf{X}}}_{i}$ $=\begin{bmatrix}\boldsymbol{x}_{t_{i,1}},\boldsymbol{x}_{t_{i,2}},\cdots,\boldsymbol{x}_{t_{i,T_{i}}}\end{bmatrix}\in\mathbb{R}^{n\times T_{i}},$
$\mathbf{U}_{i}$ $=\begin{bmatrix}\boldsymbol{u}_{t_{i,0}},\boldsymbol{u}_{t_{i,1}},\cdots,\boldsymbol{u}_{t_{i,T_{i}-1}}\end{bmatrix}\in\mathbb{R}^{m\times T_{i}},$
${\mathbf{G}}_{i}$ $=\begin{bmatrix}\boldsymbol{g}(\boldsymbol{x}_{t_{i,0}},\boldsymbol{\theta}_{i}),\cdots,\boldsymbol{g}(\boldsymbol{x}_{t_{i,T_{i}-1}},\boldsymbol{\theta}_{i})\end{bmatrix}\in\mathbb{R}^{r\times T_{i}},$
$\mathbf{\bar{G}}_{i}$ $=\begin{bmatrix}\boldsymbol{g}(\boldsymbol{x}_{t_{i,1}},\boldsymbol{\theta}_{i}),\cdots,\boldsymbol{g}(\boldsymbol{x}_{t_{i,T_{i}}},\boldsymbol{\theta}_{i})\end{bmatrix}\in\mathbb{R}^{r\times T_{i}},$
which allows (7) to be equivalently expressed as
$\mathbf{L}(A_{i},B_{i},C_{i},\boldsymbol{\theta}_{i})=\frac{1}{2T_{i}}(\parallel\bar{{\mathbf{G}}}_{i}-\begin{bmatrix}A_{i}\ B_{i}\end{bmatrix}\begin{bmatrix}{\mathbf{G}}_{i}\\ \mathbf{U}_{i}\end{bmatrix}\parallel_{F}^{2}$ (9)
$+\parallel{\mathbf{X}}_{i}-C_{i}{\mathbf{G}}_{i}\parallel_{F}^{2}).$
Multi-agent optimization problem. Based on $\mathbf{L}$ in (9), we formulate the following multi-agent optimization problem to approximate (3):
$\min_{\{A_{i},B_{i},C_{i},\boldsymbol{\theta}_{i}\}_{i=1}^{N}}$ $\sum_{i=1}^{N}\mathbf{L}(A_{i},B_{i},C_{i},\boldsymbol{\theta}_{i})$ (10)
$\text{subject to}\quad A_{1}$ $=A_{2}=\cdots=A_{N},$
$B_{1}$ $=B_{2}=\cdots=B_{N},$
$C_{1}$ $=C_{2}=\cdots=C_{N},$
$\boldsymbol{\theta}_{1}$ $=\boldsymbol{\theta}_{2}=\cdots=\boldsymbol{\theta}_{N}.$
III-B Algorithm
We now propose a distributed algorithm to solve (10), which proceeds in two steps. First, for a fixed parameter vector $\boldsymbol{\theta}_{i}$, problem (10) is minimized with respect to the dynamics matrices $A_{i}$, $B_{i}$, and $C_{i}$. Second, given the resulting fixed dynamics matrices from the first step, problem (10) is further minimized with respect to $\boldsymbol{\theta}_{i}$, as detailed below.
Step 1: Distributed Learning of Dynamics Matrices. While existing distributed optimization methods [19, 20] can be applied, they generally require all agents to adopt a common step size. To remove this restriction while preserving exponential convergence, we adapt the algorithm in [21] and derive the following distributed update scheme.
Suppose that all agents are initialized with a common parameter, i.e., $\boldsymbol{\theta}_{1}=\boldsymbol{\theta}_{2}=\cdots=\boldsymbol{\theta}_{N}$. Let $M_{i}=[A_{i},B_{i}]$, $N_{i}=[{\mathbf{G}}_{i}^{\prime},\mathbf{U}_{i}^{\prime}]^{\prime}$, and $s$ denotes the iteration index. The proposed distributed update rule for $M_{i}$ and $C_{i}$ is given as follows:
$\small{\begin{bmatrix}M_{i}^{\prime}(s+1)\\ E_{i}(s+1)\end{bmatrix}=F_{i}^{-1}\begin{bmatrix}d_{i}M_{i}^{\prime}(s)+\sum_{j\in\mathcal{N}_{i}}w_{ij}E_{j}(s)+cN_{i}\bar{{\mathbf{G}}}_{i}^{\prime}\\ -\sum_{j\in\mathcal{N}_{i}}w_{ij}M_{j}^{\prime}(s)+d_{i}E_{i}(s)\end{bmatrix}}$ (11)
and
$\small{\begin{bmatrix}C_{i}^{\prime}(s+1)\\ \hat{E}_{i}(s+1)\end{bmatrix}=\hat{F}_{i}^{-1}\begin{bmatrix}d_{i}C^{\prime}_{i}(s)+\sum_{j\in\mathcal{N}_{i}}w_{ij}\hat{E}_{j}(s)+c\mathbf{G}_{i}\mathbf{X}_{i}^{\prime}\\ -\sum_{j\in\mathcal{N}_{i}}w_{ij}C_{j}^{\prime}(s)+d_{i}\hat{E}_{i}(s),\end{bmatrix}}$, (12)
where $E_{i}\in\mathbb{R}^{r\times(r+m)}$ and $\hat{E}_{i}\in\mathbb{R}^{n\times r}$ are auxiliary matrices associated with agent $i$ initialized from arbitrary values $E_{i}(0)$ and $\hat{E}_{i}(0)$. The constant $c$ denotes a user-specified positive parameter. The weights satisfy $w_{ij}>0$ if $j\in\mathcal{N}_{i}$ and $w_{ij}=0$ otherwise, $d_{i}=\sum_{j\in\mathcal{N}_{i}}w_{ij}$, and with the additional assumption that the weights are symmetric, i.e., $w_{ij}=w_{ji}$. The matrices $F_{i}$ and $\hat{F}_{i}$ are given by
$\small F_{i}=\begin{bmatrix}d_{i}{\mathbf{I}}_{n}+cN_{i}N_{i}^{\prime}&d_{i}{\mathbf{I}}_{n}\\ -d_{i}{\mathbf{I}}_{n}&d_{i}{\mathbf{I}}_{n}\end{bmatrix},\hat{F}_{i}=\begin{bmatrix}d_{i}{\mathbf{I}}_{n}+c\mathbf{X}_{i}\mathbf{X}_{i}^{\prime}&d_{i}{\mathbf{I}}_{n}\\ -d_{i}{\mathbf{I}}_{n}&d_{i}{\mathbf{I}}_{n}\end{bmatrix}$.
By following the analysis in [21], for fixed $\boldsymbol{\theta}_{i}$, the update rule (11)-(12) converge exponentially fast to the optimal solutions of (10).
Step 2: Distributed Parameters Tuning. Suppose that $M_{i}$ and $C_{i}$ in (9) are constant matrices obtained from Step $1$, the loss function in (9) is then only determined by $\boldsymbol{\theta}_{i}$ as:
$\mathbf{L}(\boldsymbol{\theta}_{i})=\frac{1}{2T_{i}}(\parallel\bar{{\mathbf{G}}}_{i}-M_{i}\begin{bmatrix}{\mathbf{G}}_{i}\\ \mathbf{U}_{i}\end{bmatrix}\parallel_{F}^{2}+\parallel{\mathbf{X}}_{i}-C_{i}{\mathbf{G}}_{i}\parallel_{F}^{2}).$
The parameter $\boldsymbol{\theta}_{i}$ is updated following the distributed subgradient methods for multi-agent optimization in [22]:
$\boldsymbol{\theta}_{i}(s+1)=\sum_{j\in\mathcal{N}_{i}}\hat{w}_{ij}\boldsymbol{\theta}_{j}(s)-\alpha_{i}(s)\nabla_{\boldsymbol{\theta}_{i}}{\mathbf{L}}(\boldsymbol{\theta}_{i}(s)),$ (13)
where $\alpha_{i}(s)$ denotes the learning rate of agent $i$ at iteration $s$. The weights satisfy $\hat{w}_{ij}>0$ if $j\in\mathcal{N}_{i}$, and $\hat{w}_{ij}=0$ otherwise. Let $\hat{W}\in\mathbb{R}^{N\times N}$ denote the weight matrix with its $ij$-th entry $\hat{w}_{ij}$, we assume $\hat{W}$ is doubly stochastic.
In summary, we have the following algorithm referred to as a distributed deep Koopman learning using partial trajectories (DDKL-PT) in the rest of this paper.
0: Each agent observes partial trajectory $\xi_{i}$ in (4).
0: Each agent $i$ constructs $\boldsymbol{g}(\cdot,\boldsymbol{\theta}_{i}):\mathbb{R}^{n}\rightarrow\mathbb{R}^{r}$ with nonzero $\boldsymbol{\theta}_{i}(0)\in\mathbb{R}^{p}$, where $\theta_{1}(0)=\theta_{2}(0)=\cdots=\theta_{N}(0)$. Each agent $i$ initializes random matrices $M_{i}(0)\in\mathbb{R}^{r\times(r+m)}$, $C_{i}(0)\in\mathbb{R}^{n\times r}$, $E_{i}(0)\in\mathbb{R}^{r\times(r+m)}$, $\hat{E}_{i}(0)\in\mathbb{R}^{n\times r}$, sets the iteration steps $S$ and $\bar{S}$, the learning rates $\{\alpha_{i}(s)\}_{s=0}^{\bar{S}}$, and weights $w_{ij}$, $c$ and $\hat{w}_{ij}$.
1:for $t\leftarrow 0$ to $S$ do
2:  Each agent $i$ updates $M_{i}(s)$ and $C_{i}(s)$ following (11) and (12), respectively.
3:end for
4:for $t\leftarrow 0$ to $\bar{S}$ do
5:  Each agent $i$ updates $\boldsymbol{\theta}_{i}(s)$ following (13), where $\nabla_{\boldsymbol{\theta}_{i}}{\mathbf{L}}(\boldsymbol{\theta}_{i}(s))$ is computed using $M_{i}(S)$ and $C_{i}(S)$.
6:end for
7: Save the resulting $\{A_{i}(S),B_{i}(S),C_{i}(S),\boldsymbol{\theta}_{i}(\bar{S})\}$.
Algorithm 1 Distributed Deep Koopman Learning using Partial Trajectories
IV Simulations
In this section, we first evaluate the dynamics learning performance of the proposed DDKL-PT algorithm by analyzing its prediction error on a testing dataset. We then present a motivating example that demonstrates the use of DDKL-PT in model predictive control (MPC) design for a surface vehicle performing goal-tracking and station-keeping tasks for each agent in the MAS. All simulations consider a five-agent MAS, depicted in Fig. 2 (self-arcs omitted for clarity), where each agent observes only a partial trajectory $\boldsymbol{\xi}_{i}$ as defined in (4).
Figure 2: Five-agent connected network with self-arcs, where self-arcs are omitted for simplicity.
We first consider the surface vehicle dynamics from [23], characterized by a $6$-dimensional state vector $\boldsymbol{x}(t)=\begin{bmatrix}\boldsymbol{p}(t)^{\prime}&\boldsymbol{v}(t)^{\prime}\end{bmatrix}^{\prime}\in\mathbb{R}^{6}$, where $\boldsymbol{p}(t)=[p_{x}(t)\ p_{y}(t)\ \phi(t)]^{\prime}$ represents the $x$ and $y$ positions and yaw angle of the surface vehicle in global coordinates, and $\boldsymbol{v}(t)=\begin{bmatrix}v_{x}(t)&v_{y}(t)&\dot{\phi}(t)\end{bmatrix}^{\prime}$ denotes the corresponding velocities in local coordinates. The system is controlled by a $2$-dimensional control input $\boldsymbol{u}(t)=\begin{bmatrix}u_{left}&u_{right}\end{bmatrix}^{\prime}$, where $1\leq u_{left}\leq 1$ and $1\leq u_{right}\leq 1$ represent the thrusts of the left and right motors, respectively.
A nominal trajectory $\boldsymbol{\xi}$ over the time interval $0\leq t\leq 5000$ is generated by driving the surface vehicle with randomly sampled control inputs following a normal distribution. The time intervals for these partial observations of each agent are $0\leq t_{1,j}\leq 600$, $600\leq t_{2,j}\leq 1600$, $1600\leq t_{3,j}\leq 2800$, $2800\leq t_{4,j}\leq 3600$, and $3600\leq t_{5,j}\leq 4000$. The trajectory segment $4000\leq t\leq 5000$ is reserved for testing the learned models.
IV-A Dynamics Learning Evaluation
The objective is to learn the mapping from velocity $\boldsymbol{v}_{t}$ and control input $\boldsymbol{u}_{t}$ to the next velocity $\boldsymbol{v}_{t+1}$ for all agents in the MAS. For implementation of DDKL-PT, we set $w_{ij}=1$ and $c=1$ for the update rules (11)-(12). The weights $\hat{w}_{ij}$ in (13) are chosen as uniform neighbor weights, i.e., each agent assigns equal weights to its neighbors. The DNNs $\boldsymbol{g}(\boldsymbol{v}_{t},\boldsymbol{\theta}_{i}):\mathbb{R}^{3}\rightarrow\mathbb{R}^{8}$ are constructed with one hidden layer, ReLU activation, and $256$ nodes. The proposed method is compared against two baselines:
  • Centralized DKO: Learn Koopman operators by minimizing (7) over the full trajectory $0\leq t\leq 4000$, resulting in the dynamics in (3) using the same initial $\boldsymbol{\theta}$ as the proposed method.
  • Multilayer perceptron (MLP): Learns $\boldsymbol{\theta}$ by minimizing
    ${\mathbf{L}}(\boldsymbol{\theta})=\frac{1}{4000}\sum_{t=0}^{3999}\|\boldsymbol{g}(\boldsymbol{v}_{t},\boldsymbol{u}_{t},\boldsymbol{\theta})-\boldsymbol{v}_{t+1}\|^{2}.$
All methods use the same training dataset ($0\leq t\leq 4000$), testing dataset ($4000\leq t\leq 5000$), learning rate $10^{-4}$, and Adam optimizer [24]. Training is terminated once the loss drops below $7\times 10^{-6}$.
Evaluation Metrics. We measure the average prediction error over $10$ runs to mitigate the randomness of DNN training. For the proposed method, the metric is defined as
$V_{\textrm{DDKL-PT}}=\frac{1}{5\times 10^{4}}\sum_{i=1}^{5}\sum_{j=1}^{10}\sum_{t=4000}^{4999}\|\boldsymbol{\hat{v}}_{i,t+1}^{j}-\boldsymbol{v}_{t+1}\|^{2},$
where $\boldsymbol{\hat{v}}_{i,t+1}^{j}$ denotes the prediction of agent $i$ given $\boldsymbol{v}_{i,t}^{j}$ and $\boldsymbol{u}_{i,t}^{j}$ at the $j$-th run experiment. For the DKO and MLP baselines, the metric is defined as
$V=\frac{1}{10^{4}}\sum_{j=1}^{10}\sum_{t=4000}^{4999}\|\boldsymbol{\hat{v}}_{t+1}^{j}-\boldsymbol{v}_{t+1}\|^{2}.$
To mitigate the effect of randomness in DNN training, we conduct $10$ experimental runs.
Results analysis. Figs. 3(a)-3(b) illustrate that the matrices $A_{i},B_{i},C_{i}$ obtained by each agent through (11)-(12) with $w_{ij}=1$, $c=1$ converge to the dynamics matrices $A,B,C$ obtained from the centralized DKO using the complete trajectory and the same DNN $\boldsymbol{g}(\cdot,\boldsymbol{\theta})$. Similarly, the tunable parameters $\boldsymbol{\theta}_{i}$ reach consensus via the update rule (13).
(a) Consensus of $A_{i}$ and $B_{i}$ matrices.
(b) Consensus of $C_{i}$ matrix and $\boldsymbol{\theta}_{i}$.
Figure 3: Consensus of the DDKR of each agent $i$ during the learning process, where the solid black line denotes the average of the 1010 experiments, and the blue shadow denotes the standard deviation.
A one-way analysis of variance (ANOVA) [25, Chapter 14] was performed to compare testing errors among the three methods: $V_{\textrm{DDKL-PT}}$ (proposed, distributed), $V_{\textrm{DKO}}$ (centralized), and $V_{\textrm{MLP}}$ (centralized). The results reveal a statistically significant difference in errors between at least two groups, with a p-value of $0.00<0.05$. Tukey’s HSD test further establishes the ranking DKO << MLP << DDKL-PT. Table I summarizes the testing errors. The slightly higher error of the proposed DDKL-PT method arises from the distributed nature of the dataset across agents, in contrast to the centralized training of DKO and MLP. Nevertheless, the simulations in Section IV-B demonstrate that the Koopman representations obtained by DDKL-PT provide sufficient accuracy for downstream optimal control tasks.
Methods Estimation errors on testing dataset
[…]
TABLE I: Average estimation errors and standard deviations computed over the testing data across 1010 experimental runs.
IV-B MPC Design with Learned DDKL-PT Dynamics
Consider any agent $i$ in Fig. 2. An MPC controller is designed using the learned dynamics obtained from the proposed DDKL-PT framework to drive the surface vehicle from the initial condition $x(0)=\begin{bmatrix}20&10&\pi/3&0&0&0\end{bmatrix}^{\prime}$ to goal state $x_{goal}=\begin{bmatrix}0&0&\pi/2&0&0&0\end{bmatrix}^{\prime}$.
Combined Dynamics of DKO and Kinematics: The learned deep Koopman dynamics are expressed as
$\boldsymbol{v}(t+1)=C_{i}^{*}\Big(A_{i}^{*}\boldsymbol{g}(\boldsymbol{v}(t),\boldsymbol{\theta}_{i}^{*})+B_{i}^{*}\boldsymbol{u}(t)\Big)=\begin{bmatrix}v_{x}(t+1)\\ v_{y}(t+1)\\ \dot{\phi}(t+1)\end{bmatrix}$, (14)
where $A_{i}^{*}$, $B_{i}^{*}$, $C_{i}^{*}$, and $\boldsymbol{\theta}_{i}^{*}$ come from Section IV-A. This model is employed to predict the future velocities of the surface vehicle in local coordinates.
To obtain full-state predictions, the learned Koopman model (14) is integrated with the vehicle’s kinematic dynamics:
$\boldsymbol{p}(t+1)=\boldsymbol{p}(t)+\boldsymbol{\dot{p}}(t)\Delta t,\quad\Delta t=0.02s,$ (15)
where $\boldsymbol{\dot{p}}(t)$ is computed based on the DKO-predicted velocities from (14) as
$\boldsymbol{\dot{p}}(t)=\begin{bmatrix}\cos(\phi(t))v_{x}(t+1)-\sin(\phi(t))v_{y}(t+1)\\ \sin(\phi(t))v_{x}(t+1)+\cos(\phi(t))v_{y}(t+1)\\ \dot{\phi}(t+1)\end{bmatrix}$.
The combined dynamics are therefore defined as
$\boldsymbol{x}(t+1)=\boldsymbol{F}(\boldsymbol{x}(t),\boldsymbol{u}(t))=\begin{bmatrix}\boldsymbol{p}(t+1)\\ \boldsymbol{v}(t+1)\end{bmatrix}$, (16)
where $\boldsymbol{p}(t+1)$ and $\boldsymbol{v}(t+1)$ are given in (15) and (14), respectively.
MPC Setup. At each time step $t=0,1,2,\cdots$ the following MPC problem is solved for agent $i$:
[…]
with horizon length $K=30$. The stage cost is defined as
$c_{t}=(\boldsymbol{x}(t)-\boldsymbol{x}_{goal})^{\prime}Q(\boldsymbol{x}(t)-\boldsymbol{x}_{goal})+\boldsymbol{u}(t)^{\prime}R\boldsymbol{u}(t),$
and the terminal cost as
$c_{f}=(\boldsymbol{x}(K)-\boldsymbol{x}_{goal})^{\prime}Q_{f}(\boldsymbol{x}(K)-\boldsymbol{x}_{goal})$
where $Q=\text{diag}(300,300,500,10,10,10)$, $Q_{f}=2Q$ and $R=\text{diag}(10^{-3},10^{-3})$ are chosen positive definite diagonal weight matrices.
Results Analysis. Simulation results are shown in Fig. 4.
Figure 4: Tracking errors of each agent in MAS driven by DDKL-PT-MPC
By applying MPC with the deep Koopman model learned via DDKL-PT, each agent $i$ successfully reaches the goal state. The convergence time is approximately $300$ time steps across agents..
V Conclusions
This paper presented a distributed deep Koopman learning algorithm (DDKL-PT) for approximating the dynamics of nonlinear time-invariant systems in a multi-agent setting. Each agent learns the unknown dynamics from partial trajectory segments and exchanges its locally estimated dynamics with neighbors to achieve consensus, thereby preserving data privacy without sharing raw training data. By integrating the learned Koopman dynamics with a kinematic model, a model predictive control (MPC) framework was developed. Simulation results on a surface vehicle demonstrated both the accuracy of the distributedly learned dynamics and the effectiveness of the proposed MPC for goal-tracking and station-keeping tasks. All agents reached the desired goal state within approximately $300$ time steps, confirming the capability of DDKL-PT for distributed learning and model-based optimal control.